OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

jhonyy9 (jhonyy9):

how do you think it ?

jhonyy9 (jhonyy9):

so like first step subtract from both sides 3x so than will get ?

jhonyy9 (jhonyy9):

3x-3x-7=5x-3x+19 0 -7=2x+19 add to both sides -19 so will get -7-19=2x+19-19 -26=2x+0 divide both sides by 2 -26/2 =2x/2 -13=x ok ?

jhonyy9 (jhonyy9):

3*(-13)-7=5(-13)+19 -39-7=-65+19 -46=-46

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

first, move all terms that have x into same side, others into the other so you will get 2x = -26 then you divide both sides by 2, and you will get your answer